you have been given the mathematical model to calculate velocity of a car accelerating from rest in straight line: v(t) = A (1-e ^ - t/tmaxspeed) (anything after the symbol ^ is to the power of) v(t is instantaneous velocity of car (m/s), t is time in seconds, tmaxspeed is time to reach maximum speed in seconds and A is constant. you are given the information that t (0-28 m/s) is 2.6s, t (400m) is 10.46s & tmaxspeed is 7s. please explain and derive an equation x(t) for the instantaneous position of the car as a function of time.....identify the value of x when t=0s ? and asymptote of this function as t → ∞?
1
Expert's answer
2020-02-26T13:17:00-0500

Given : "v(t) = A (1-e ^{-t\/t_m})" , where "t_m" is time to reach maximum speed in seconds and A is constant.

Also given; "v(2.6)=28m\/s;t_m=7s"

"\implies v(t) = A (1-e ^{-t/7})"

"\implies v(2.6) = A (1-e ^{-2.6/7})=28"

"\implies A=90.25m/s"

"\implies v(t) =90.25(1-e ^{-t/7})=dx/dt"

Integrating on both sides, we get;

"\implies \int dx=90.25 \int (1-e^{-t/7})dt"

"\implies x(t)=90.25(t+e^{-t/7}/7)+c"

Also given; "x(10.46)=400=90.25(10.46+e^{-10.46\/7}\/7)+c=964.268+c"

"\implies c=-564.268 m"

"\implies x(t)=90.25(t+e^{-t/7}/7)-564.268" ---(1)

(1) gives instantaneous position of the car as a function of time.

"x(t=0)=90.25(0+e^0\/7)-564.268=90.25\/7-564.268=-551.375"

"\implies x(0)=-551.375m" ----(Answer)

as "t \to \infty, e^{-t} \to 0 \implies x \to \infty"

"k=\lim_{t \to \infty} \frac{x(t)}{t}=90.25," "b=\lim_{t \to \infty}(x(t)-kt)=\lim_{t \to \infty}(x(t)-90.25t)=-564.268"

The asymptote of this function is "y(t)=90.25t-564.268".